OpenStudy (anonymous):

Solve the equation. If necessary, round your answer to the nearest tenth. x^2=196 A. 14 B. -14,14 C. 98 D. -98,98 Help please? (: will give a medal.

OpenStudy (anonymous):

c

OpenStudy (jdoe0001):

\(\bf x^2=196 \implies \sqrt{x^2}=\sqrt{196}\implies x=\pm \sqrt{196}\)